Free CFA I Exam Braindumps
CFA Level I Chartered Financial Analyst Questions 1086–1090 (page: 219)

Updated On: 9-Feb-2026

If you deposit $1,500 into an account paying 6% per year, compounded quarterly, how much is in the account after 60 months?

  1. $2,224.46
  2. $2,117.54
  3. $2,020.28
  4. $1,702.48
  5. $1,751.59

Answer(s): C

Explanation:

There are 20 quarters in 60 months (60 divide 3 =). On the BAII Plus, press 20 N, 6 divide 4 = I/Y, 1500 PV, 0 PMT, CPT FV. On the HP12C, press 20 n, 6 ENTER 4 divide i, 1500 PV, 0 PMT, FV. Note that the answer will be displayed as a negative number. Make sure the BAII Plus has the value of P/Y set to 1.



BeigeBox, a computer manufacturer, can construct computer systems with any one of 4 processors, 2 memory configurations, 3 hard drives, 2 monitors, 2 keyboards, and 2 CD drives. How many unique configurations of computer systems can BeigeBox construct?

  1. 192.
  2. 96.
  3. 384.
  4. 64.

Answer(s): A

Explanation:

The multiplication rule of counting states that the number of combinations available when there are n_1 options in one aspect, n_2 in another, and so on, up to n_k, is n_1
* n_2 * ... * n_k. In this case, the number of combinations is 4*2*3*2*2*2 = 192.



In a beer taste study involving the ExtraFerm lager beer, the following responses were recorded:
Response Frequency
Great taste 103
Okay taste 46
Bad taste 64
The probability that a randomly selected beer drinker will NOT find ExtraFerm to be bad tasting is calculated from these data to be:

  1. 0.22
  2. 0.30
  3. 0.70
  4. 0.78

Answer(s): C

Explanation:

The probability that a randomly selected beer drinker will find ExtraFerm to be bad tasting is estimated from the above data as 64/(64 + 46 + 103) = 0.3. Therefore, the probability that a randomly selected beer drinker will NOT find ExtraFerm to be bad tasting equals 1 - 0.3 = 0.7. We could also have calculated this probability as (103 + 46)/(64 + 46 + 103) = 0.7 Note that given a frequency distribution, the estimated probability of a particular class/event occurring equals the relative frequency of that class.



A derivative security pays $100 if the Dow Jones Industrial Average shows a 15% return over the next year and does not pay anything of the return is lower. If the security costs $45 and its expected return is 30%, what's the probability that the Dow Jones return will exceed 15%?

  1. 39.87%
  2. 58.51%
  3. 41.49%
  4. 61.22%

Answer(s): B

Explanation:

If p is the probability of the Dow Jones return exceeding 15%, then the expected total return on the derivative security is (p * 100 + (1-p)*0)/45 = 2.222p. Since the expected return is given to be 30%, we get 2.22p = 1.3.
Therefore, p = 1.3/2.222 = 58.51%



A bottling company offers three kinds of delivery services-instant, same day and within five days. The profit per delivery varies according to the kind of delivery. The profit for an instant delivery is less than the other kinds because the driver has to go directly to a grocery store with a small load and return to the bottling plan. To find out what effect each type of delivery has on the profit picture, the company has made the following tabulation based on deliveries for the previous quarter.

Type of Number of Deliveries Profit per Delivery
Delivery During the Quarter
Instant 100 $70
Same day 60 $100
Within five days 40 $160

What is the weighted mean profit per delivery?

  1. $97
  2. $72
  3. $142
  4. None of these answers
  5. $100

Answer(s): A

Explanation:

(100*70)+(60*100)+(40*160) = 1940. Mean is 1940/200 = 97



Viewing page 219 of 793
Viewing questions 1091 - 1095 out of 3960 questions


Post your Comments and Discuss CFA CFA I exam prep with other Community members:

AI Tutor AI Tutor 👋 I’m here to help!